Session: May 2002 (session 2)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Tonino Pisanu at Sat May 25 07:34:44 CEST 2002) Success
Any problem

Noto (P.Cassaro at Sat May 25 15:57:24 CEST 2002) Success
Scan lost at 00:50 for prelimit.

Effelsberg (A. Kraus at Mon May 27 09:34:45 CEST 2002) Success

Westerbork (Tony Foley at Mon May 27 12:25:13 CEST 2002) Success
No known problems

Onsala (John Conway at Tue May 28 09:27:45 CEST 2002) Success
No known problems

Jodrell_Bank (Alastair Gunn at Tue May 28 14:56:55 CEST 2002) Minor failures
Telescope offline due to high winds until start of second tape at 23:13 UT. Only one tape sent to correlator. No phase CAL present. No other known problems.